英語 (English)

発信型英語 (Communicative English) (総科 ((人文 E)2 年))

ポンドクリストファー · 非常勤講師 / 全学共通教育センター 2 単位前期月 1 ・ 2

(平成 19 年度以前の授業科目: 『発信型英語』 ) (平成 16 年度以前 (医保は 17 年度以前) の授業科目: 『英語 (2)』 )

【授業の目的】 In today’s globalized society the ability to communicate effectively in English is becoming more and more important. This class helps students to develop their communicative ability through the use of in class communicative activities around various, everyday topics. Students will also be given the opportunity to improve other skills such as reading ability, grammatical knowledge, listening skills and as well as build their level of vocabulary.

【授業の概要】 This is a communicative style class. Students will be expected to take part in various pair-work and small group activities such as role-plays, discussions and information-gap tasks. There will be a short homework activity each week that is necessary preparation for the following week’s class. The mid-term and final exams are presentation based.

【キーワード】 [キーワード] 【先行科目】 [先行科目] 【関連科目】 [関連科目] 【到達目標】 By the end of this course students will have gained the communicative

ability, vocabulary, listening skills and grammar necessary to take part in discussions on everyday topics.

【授業の計画】 1. Introduction to the course, course outline and requirements.Introductions

2. Getting to know you talking about you, friends, family and neighbors.Tenses and questions

3. Introductory presentations

4. Getting to know you, talking about friends, family and neighbors. Listening and speaking

5. Getting to know you, talking about friends, family and neighbors. Reading and speaking

6. Everyday English and social expressions. Using a bilingual dictionary.

7. The way we live. Present tenses

8. Mid-term presentation

9. The way we live. Vocabulary and speaking 10. The tales of two cities. Reading and speaking

11. 24/7 society listening and speaking, everyday English

12. Famous people. Present perfect, past simple, for and since

13. Music. Listening and speaking

14. Writing a biography

15. Final presentation

16. Review and feedback

【教科書】 New Headway 3rd Edition Pre-intermediate STUDENT BOOK, Liz and

John Soars, Oxford University Press, ISBN-13: 9780 1947 15850 【参考書等】 [参考資料] 【成績評価の方法】 Weekly attendance, active participation and homework are an important part of this course. Students are evaluated as follows: Attendance, participation and homework 50%, mid term presentation 15%, final presentation 35%

【再試験の有無】 Allowable if the student has attended more than 2/3 of the classes

and has a score of 50% or higher. 【受講者へのメッセージ】 [注意]

【授業コンテンツ】 http://cms.db.tokushima-u.ac.jp/cgi-bin/toURL?EID=220885 【連絡先 (オフィスアワー・研究室・ E メールアドレス)】 ⇒ ポンド .

English

Communicative English (総科 ((人文 E)2 年))

Christopher Pond · Part-time Lecturer / Center for General Education 2 units 前期月 1 ・ 2

(平成 19 年度以前の授業科目: 『発信型英語』 ) (平成 16 年度以前 (医保は 17 年度以前) の授業科目: 『英語 (2)』 )

Target⟩ In today’s globalized society the ability to communicate effectively in English is becoming more and more important. This class helps students to develop their communicative ability through the use of in class communicative activities around various, everyday topics. Students will also be given the opportunity to improve other skills such as reading ability, grammatical knowledge, listening skills and as well as build their level of vocabulary.

Outline⟩ This is a communicative style class. Students will be expected to take part in various pair-work and small group activities such as role-plays, discussions and information-gap tasks. There will be a short homework activity each week that is necessary preparation for the following week’s class. The mid-term and final exams are presentation based.

�������⟩ [キーワード] Fundamental Lecture⟩ [先行科目] Relational Lecture⟩ [関連科目] Goal⟩ By the end of this course students will have gained the communicative

ability, vocabulary, listening skills and grammar necessary to take part in discussions on everyday topics.

Schedule⟩

1. Introduction to the course, course outline and requirements.Introductions

2. Getting to know you talking about you, friends, family and neighbors.Tenses and questions

3. Introductory presentations

4. Getting to know you, talking about friends, family and neighbors. Listening and speaking

5. Getting to know you, talking about friends, family and neighbors. Reading and speaking

6. Everyday English and social expressions. Using a bilingual dictionary.

7. The way we live. Present tenses

8. Mid-term presentation

9. The way we live. Vocabulary and speaking 10. The tales of two cities. Reading and speaking

11. 24/7 society listening and speaking, everyday English

12. Famous people. Present perfect, past simple, for and since

13. Music. Listening and speaking

14. Writing a biography

15. Final presentation

16. Review and feedback

Textbook⟩ New Headway 3rd Edition Pre-intermediate STUDENT BOOK, Liz

and John Soars, Oxford University Press, ISBN-13: 9780 1947 15850 Reference⟩ [参考資料] Evaluation Criteria⟩ Weekly attendance, active participation and homework are an important part of this course. Students are evaluated as follows: Attendance, participation and homework 50%, mid term presentation 15%, final presentation 35%

Re-evaluation⟩ Allowable if the student has attended more than 2/3 of the classes

and has a score of 50% or higher. Message⟩ [注意]

Contents⟩ http://cms.db.tokushima-u.ac.jp/cgi-bin/toURL?EID=220885 Contact (Office-Hour, Room, E-mail)⟩

⇒ Pond .